### v1

@@ Abstract
Large Language Models (LLMs) and multi-agent systems have shown impressive capabilities in natural language tasks but face challenges in clinical trial applications, primarily due to limited access to external knowledge. Recognizing the potential of advanced clinical trial tools that aggregate and predict based on the latest medical data, we propose an integrated solution to enhance their accessibility and utility. We introduce Clinical Agent System (CT-Agent), a Clinical multi-agent system designed for clinical trial tasks, leveraging GPT-4, multi-agent architectures, LEAST-TO-MOST, and ReAct reasoning technology. This integration not only boosts LLM performance in clinical contexts but also introduces novel functionalities. Our system autonomously manages the entire clinical trial process, demonstrating significant efficiency improvements in our evaluations, which include both computational benchmarks and expert feedback.
1 Introduction
@@
The introduction of clinical multi-agent systems into the healthcare sector marks a substantial advancement in improving care quality through sophisticated computational methods and in-depth data analysis. These systems, driven by Large Language Models (Singhal et al., 2023a) (LLMs) like ChatGPT (Liu et al., 2023), BioGPT (Luo et al., 2022), ChatDoctor (Yunxiang et al., 2023), and Med-PaLM (Singhal et al., 2023b), have shown considerable success in processing and understanding medical data, providing customized care, and offering insights into intricate health conditions. However, their use in clinical trials faces challenges, mainly due to their limited ability to access and integrate external knowledge sources, such as DrugBank (Wishart et al., 2018). This research stems from the urgent need to fully utilize LLMs in clinical settings, going beyond the conversational skills of current models to include actionable and explanatory analysis leveraging extensive external data.
@@
Our study introduces CT-Agent, a new Clinical multi-agent system tailored for clinical trial tasks. Utilizing the capabilities of GPT-4, combined with multi-agent system architectures, and incorporating advanced reasoning technologies like LEAST-TO-MOST (Zhou et al., 2022) and ReAct (Yao et al., 2022), our solution not only boosts LLM performance in clinical scenarios but also brings new functionalities. Our system is designed to autonomously oversee the clinical trial process, filling the void in existing implementations that mainly focus on conversational interactions without sufficient actionable outcomes.
@@
Prior research has highlighted the potential of LLMs in healthcare, particularly in diagnostics, patient communication, and medical research (Singhal et al., 2023b; Yunxiang et al., 2023; Singhal et al., 2023a). Yet, these investigations have not fully exploited the models for clinical trials, where understanding the complex relationships between drugs, diseases, and patient reactions is crucial. Our research introduces a multi-agent framework that uses specialized agents for tasks such as drug information retrieval, disease analysis, and explanatory reasoning. This strategy not only allows for a more detailed and understandable decision-making process but also significantly enhances clinical trial analysis capabilities, including predicting outcomes, deciphering reasons for failure, and estimating trial duration.
@@
We present Clinical Trial Multi-Agent (CT-Agent), the first multi-agent framework that elevates the conversational abilities of LLMs with actionable intelligence.
@@
CT-Agent achieves competitive predictive performance in clinical trial outcome prediction (0.7908 PR-AUC), obtaining a 0.3326 improvement over the Standard Prompt Method.
2 Related Work
NLP has achieved significant progress in the biomedical arena, delivering crucial insights and tools for a range of applications in healthcare and medicine. The advent of Large Language Models (LLMs) has notably advanced the medical field by embedding comprehensive medical knowledge into their training.
For example, question answering (QA) in the medical domain represents a critical challenge in NLP, where language models are tasked with responding to specific queries using their embedded medical knowledge, e.g., MedQA (USMLE) (Jin et al., 2020) HeadQA (Vilares and Gómez-Rodríguez, 2019), MMLU (Hendrycks et al., 2021), and PubMedQA (Jin et al., 2019).
Despite being pretrained for general purposes, closed-source LLMs like ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023) have demonstrated considerable medical capabilities in both benchmark evaluations and real-world applications. Liévin et al. (2023) applied GPT-3.5 using various prompting techniques, such as Chain-of-Thought, few-shot, and retrieval augmentation, across three medical reasoning benchmarks, showcasing the model’s robust medical reasoning skills without the need for specialized fine-tuning. Additionally, evaluations of LLMs such as ChatGPT on professional medical assessments, including the US Medical Exam (Kung et al., 2023) and the Otolaryngology-Head and Neck Surgery Certification Examinations (Long et al., 2023), have resulted in scores that meet or nearly meet passing thresholds. This performance underscores the potential of LLMs to aid in significant medical contexts, including medical education and clinical decision-making.
@@ 2 Related Work
AI has great potential to revolutionize clinical trials in a couple of problems. Specifically, Zhang et al. (2020); Gao et al. (2020) leverage AI to recruit appropriate patients that meet the requirement in eligibility criteria. Fu et al. (2022); Chen et al. (2024); Lu et al. (2024) builds machine learning models to predict the outcome of clinical trials based on clinical trial features such as drug molecule, disease code, and eligibility criteria. In the context of clinical trials, Wang et al. (2024) leverages the large language model to generate patient-level digital twins to simulate clinical trials. However, most of these works do not utilize LLM’s reasoning ability and cannot solve complex problems in clinical trials. To the best of our knowledge, our work is the first LLM agent work for clinical trials and is able to solve complex clinical trial reasoning problems.
3 Methods
3.1 Overview of CT-Agent
@@
Our proposed system is a conversational multi-agent framework, analogous to a hospital staffed by various specialists. Each agent within this system plays a distinct role, mirroring the specialization seen in medical professionals—some focus on pharmacology, others on diagnosing diseases, while a few are dedicated to designing clinical trials. To process natural language inputs and generate responses that are coherent and contextually appropriate, each agent utilizes GPT-4. Moreover, we enhance the system’s reasoning capabilities by incorporating methodologies such as ReAct (Yao et al., 2022) and the LEAST-TO-MOST (Zhou et al., 2022) principle. Following the reasoning process, the system is capable of taking actions such as searching for information, indexing data in databases, and employing expert AI models. By integrating this information, the system effectively simulates a highly knowledgeable doctor. Working in concert, these agents can deliver precise, explainable solutions to user inquiries.
3.2 Agent Roles and Responsibilities
@@
The CT-Agent framework integrates a diverse array of specialized agents, each employing the ReAct and LEAST-TO-MOST reasoning methods to meticulously plan their actions. Through the use of advanced search capabilities, access to specialist models, and indexing in databases, these agents are able to execute a wide range of tasks effectively. Below, we delve into the specific roles and responsibilities assigned to each agent within the system.
3.2.1 Planning Agent
@@
3.2.2 Efficacy Agent
The Efficacy Agent is a specialized module within our multi-agent framework, primarily focused on assessing the therapeutic effectiveness of drugs against specified diseases (Chang et al., 2019; Chen et al., 2021; Zhang et al., 2021; Lu et al., 2022). This agent utilizes advanced data retrieval and analysis techniques, drawing from rich biomedical databases such as DrugBank (Wishart et al., 2018) and the HetioNet Knowledge Graph to ensure comprehensive and accurate evaluations.
@@ 3.2.2 Efficacy Agent
3.2.3 Safety Agent
The Safety Agent is integral to our CT-Agent framework, focusing specifically on the assessment of drug safety and its implications for patient health. This agent leverages a comprehensive repository of pharmacological data and historical clinical trial outcomes to evaluate the risks associated with specific drug-disease interactions. Utilizing databases such as DrugBank and clinical trial registries, the Safety Agent provides detailed insights into the historical safety profiles of drugs.
@@ 3.2.3 Safety Agent
3.2.4 Enrollment Agent
Proper enrollment ensures that the trial has enough participants to statistically power the study. This is essential to detect the true effect of the intervention being tested. Insufficient enrollment can lead to inconclusive or unreliable results because the sample size determines the ability of a trial to accurately reflect the effects of a treatment. We leverage a Hierarchical transformer-based model Yue et al. (2024) that takes eligibility criteria as an input feature and predicts the success rate of enrollment. It is a binary classification problem, where 1 denotes the successful enrollment while 0 does not. The details can be found in Section 3.3.
3.3 Calling External Tools
@@
Drugbank: Drugbank (Wishart et al., 2018) stands out as a premier resource, offering detailed drug data, including chemical, pharmacological, and pharmaceutical information, with a focus on comprehensive drug-target interactions. Drugbank is not only a repository of drug information but also serves as an invaluable tool for bioinformatics and cheminformatics research, providing data for over 13,000 drug entries including FDA-approved small molecule drugs, FDA-approved biopharmaceuticals (proteins, peptides, vaccines, and allergenics), and nutraceuticals.
Hetionet: Hetionet (Himmelstein et al., 2017) is an integrative network of biology that encompasses a comprehensive collection of biological entities and their relationships. It uniquely combines data from various biomedical databases covering diseases, genes, compounds, and more, into a single, coherent graph structure. This interconnected approach allows for multifaceted analyses, including drug repurposing, genetic associations, and network medicine. Hetionet includes over 47,000 nodes of different types (e.g., diseases, drugs, genes) and more than 2 million relationships, offering a rich dataset for computational biology and drug discovery.
@@
Enrollment Model: The enrollment model is designed to predict the likelihood of successful participant enrollment in clinical trials based on the eligibility criteria, the drugs involved, and the diseases targeted. This is a hierarchical transformer-based model, integrating sentence embeddings from BioBERT (Lee et al., 2020) to capture the nuanced medical semantics in the criteria text. In practice, the Enrollment Agent receives a query containing the drugs, diseases, and detailed eligibility criteria. It processes this information to predict the enrollment difficulty, which aids in planning and adjusting recruitment strategies for clinical trials. This capability supports more efficient trial design and can significantly impact the speed and success of new drug developments. The details can be found in A.1.
@@
3.4 Integration of Reasoning Technology
To further enhance the agent’s decision-making capabilities, we integrate advanced reasoning technologies such as ReAct (recognition, action, and context) (Yao et al., 2022) and the Least-to-Most reasoning framework (Zhou et al., 2022). These methodologies complement each other by providing robust mechanisms for addressing complex problems through structured and contextual analysis.
@@ 3.4 Integration of Reasoning Technology
3.5 Workflow
The workflow of our CT-Agent system is designed to optimize the collaboration and efficiency of multiple specialized agents to address complex medical inquiries. The process is structured in several sequential steps, as described below:
@@ 3.5 Workflow
This structured workflow ensures that CT-Agent effectively mimics a collaborative team of medical specialists, offering precise and comprehensive solutions to complex medical inquiries.
4 Experiment
This section outlines the experimental design used to assess the performance of CT-Agent in the setting of clinical trials. We aim to demonstrate the superior predictive capabilities of our model by comparing it with established baseline methods.
4.1 Baseline Methods
@@
This comparative analysis will help in highlighting the strengths and potential areas for improvement in CT-Agent, guiding future enhancements in the model’s architecture and its application in clinical trial settings.
4.2 Experimental Setup
Our experimental framework was implemented on a server equipped with an AMD Ryzen 9 3950X CPU, 64GB RAM, and an NVIDIA RTX 3080 Ti GPU. We utilized Python 3.8 for scripting and PyTorch for model implementation and training. For each experiment, we used the same seed to ensure reproducibility.
4.2.1 Data and Resources
In our investigation, we employed a variety of datasets and external resources to construct a comprehensive experimental framework. These are detailed as follows:
Drug Databases: We used DrugBank, a comprehensive, freely accessible online database containing information on drugs and drug targets. DrugBank is instrumental for acquiring detailed pharmacological information, which supports the pharmacokinetics and molecular mechanisms of drug actions in our models.
Knowledge Graphs: Hetionet, an integrative knowledge graph of biomedical information that interlinks biological entities through relationships, was utilized. It provides a structured form of data that helps in understanding complex drug-disease relationships, gene interactions, and more, which are crucial for our predictive analytics.
GPT-Generated Data: To augment our dataset and test the robustness of our models under varied scenarios, we generated synthetic data. This was done by prompting Generative Pre-trained Transformer (GPT) models with scenarios derived from our existing datasets, thereby creating realistic, hypothetical data scenarios for further testing.
These diverse resources were meticulously integrated into our model, referred to hereafter as CT-Agent. This integration helps in providing a rich and informed context for each predictive task undertaken by our model.
For our experimental validation, we randomly selected 40 training samples from the clinical trial outcome prediction benchmark provided in (Fu et al., 2022). An identical approach was used to select 40 samples from the test set, ensuring that our training and testing datasets were balanced and randomized, thereby providing a robust evaluation framework for our model’s performance.

### v2

@@ Abstract.
Large Language Models (LLMs) and multi-agent systems have shown impressive capabilities in natural language tasks but face challenges in clinical trial applications, primarily due to limited access to external knowledge. Recognizing the potential of advanced clinical trial tools that aggregate and predict based on the latest medical data, we propose an integrated solution to enhance their accessibility and utility. We introduce Clinical Agent System (ClinicalAgent), a clinical multi-agent system designed for clinical trial tasks, leveraging GPT-4, multi-agent architectures, LEAST-TO-MOST, and ReAct reasoning technology. This integration not only boosts LLM performance in clinical contexts but also introduces novel functionalities. The proposed method achieves competitive predictive performance in clinical trial outcome prediction (0.7908 PR-AUC), obtaining a 0.3326 improvement over the standard prompt Method. Publicly available code can be found at https://anonymous.4open.science/r/ClinicalAgent-6671.
1. Introduction
@@
The introduction of clinical multi-agent systems into the healthcare sector marks a substantial advancement in improving care quality through sophisticated computational methods and in-depth data analysis. Modern medicine increasingly relies on advanced technologies to enhance patient outcomes, streamline clinical processes, and provide deeper insights into complex health conditions. These advancements are driven by the integration of Large Language Models (Singhal et al., 2023a) (LLMs) like ChatGPT (Liu et al., 2023), BioGPT (Luo et al., 2022), ChatDoctor (Yunxiang et al., 2023), and Med-PaLM (Singhal et al., 2023b), which have shown considerable success in processing and understanding medical data, providing customized care, and offering insights into intricate health conditions. However, their use in clinical trials faces challenges, mainly due to their limited ability to access and integrate external knowledge sources, such as DrugBank (Wishart et al., 2018). This research stems from the urgent need to fully utilize LLMs in clinical settings, going beyond the conversational skills of current models to include actionable and explanatory analysis leveraging extensive external data.
@@
Our study introduces ClinicalAgent, a new Clinical multi-agent system tailored for clinical trial tasks. Utilizing the capabilities of GPT-4, combined with multi-agent system architectures, and incorporating advanced reasoning technologies like LEAST-TO-MOST (Zhou et al., 2023) and ReAct (Yao et al., 2023), our solution not only boosts LLM performance in clinical scenarios but also brings new functionalities. Our system is designed to autonomously oversee the clinical trial process, filling the void in existing implementations that mainly focus on conversational interactions without sufficient actionable outcomes.
@@
Prior research has highlighted the potential of LLMs in healthcare, particularly in diagnostics, patient communication, and medical research (Singhal et al., 2023b; Yunxiang et al., 2023; Singhal et al., 2023a). Yet, these investigations have not fully exploited the models for clinical trials, where understanding the complex relationships between drugs, diseases, and patient reactions is crucial (Chang et al., 2019; Lu et al., 2023b). Our research introduces a multi-agent framework that uses specialized agents for tasks such as drug information retrieval, disease analysis, and explanatory reasoning. This strategy not only allows for a more detailed and understandable decision-making process but also significantly enhances clinical trial analysis capabilities, including predicting outcomes, deciphering reasons for failure (Chen et al., 2024a), and estimating trial duration (Yue et al., 2024a).
@@
We present Clinical Trial Multi-Agent System (ClinicalAgent), the first multi-agent framework that elevates the conversational abilities of LLMs with actionable intelligence.
@@
ClinicalAgent achieves competitive predictive performance in clinical trial outcome prediction (0.7908 PR-AUC), obtaining a 0.3326 improvement over the standard large language model using prompt.
In the following sections, we discuss the related works, explore the methodology, detail the structure and functionality of our multi-agent system, outline our experimental strategy, and discuss the significance of our findings within the clinical trial context.
2. Related Work
Natural Language Processing (NLP) has achieved significant progress in the biomedical arena, delivering crucial insights and tools for a range of applications in healthcare and medicine. The advent of Large Language Models (LLMs) has notably advanced the medical field by embedding comprehensive medical knowledge into their training.
For example, question answering (QA) in the medical domain represents a critical challenge in NLP, where language models are tasked with responding to specific queries using their embedded medical knowledge, e.g., MedQA (USMLE) (Jin et al., 2021) HeadQA (Vilares and Gómez-Rodríguez, 2019), MMLU (Hendrycks et al., 2021), and PubMedQA (Jin et al., 2019).
Despite being pre-trained for general purposes, closed-source LLMs like ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023) have demonstrated considerable medical capabilities in both benchmark evaluations and real-world applications. (Liévin et al., 2023) applied GPT-3.5 using various prompting techniques, such as Chain-of-Thought, few-shot, and retrieval augmentation, across three medical reasoning benchmarks, showcasing the model’s robust medical reasoning skills without the need for specialized fine-tuning. Additionally, evaluations of LLMs such as ChatGPT on professional medical assessments, including the US Medical Exam (Kung et al., 2023) and the Otolaryngology-Head and Neck Surgery Certification Examinations (Long et al., 2023), have resulted in scores that meet or nearly meet passing thresholds. This performance underscores the potential of LLMs to aid in significant medical contexts, including medical education and clinical decision-making.
@@ 2. Related Work
AI has great potential to revolutionize clinical trials and, more generally, the biomedical industry in a couple of problems. Specifically, TrialBench (Chen et al., 2024b) highlights eight critical issues in clinical trials that are ready for AI solutions, including the forecast of trial duration, estimation of patient dropout rates, prediction of serious adverse events, mortality rates, outcomes of trial approvals, identification of trial failure reason, optimization of drug dosing, and the design of eligibility criteria. (Zhang et al., 2020; Gao et al., 2020) leverage AI to recruit appropriate patients that meet the requirement in eligibility criteria. (Fu et al., 2022, 2023; Lu et al., 2024) builds machine learning models to predict the outcome of clinical trials based on clinical trial features such as drug molecule, disease code, and eligibility criteria. To predict clinical trial enrollment success, (Yue et al., 2024b) leverages large language model-augmented features and customized deep & cross network to model the text feature. Multi-omics data enables finer-grained analysis, which is crucial for precision medicine approaches (personalized therapy). By understanding the genetic (Lu et al., 2021, 2022; Zhang et al., 2021), transcriptomic (Lu et al., 2023a; Fu et al., 2024), and other molecular profiles of patients (Chen et al., 2021), treatments can be customized to match individual disease mechanisms/biological pathway (Wu et al., 2022), potentially leading to more effective and personalized therapies (Yi et al., 2018). In the context of clinical trials, Wang et al. (2024) leverages the large language model to generate patient-level digital twins to simulate clinical trials. However, most of these works do not utilize LLM’s reasoning ability and cannot solve complex problems in clinical trials. To the best of our knowledge, our work is the first LLM agent work for clinical trials and is able to solve complex clinical trial reasoning problems.
3. Methods
3.1. Overview of ClinicalAgent
@@
Our proposed system is a conversational multi-agent framework analogous to a hospital staffed by various specialists. Each agent within this system plays a distinct role, mirroring the specialization seen in medical professionals—some focus on pharmacology, others on diagnosing diseases, while a few are dedicated to designing clinical trials. To process natural language inputs and generate responses that are coherent and contextually appropriate, each agent utilizes GPT-4. Moreover, we enhance the system’s reasoning capabilities by incorporating methodologies such as ReAct (Yao et al., 2023) and the LEAST-TO-MOST (Zhou et al., 2023) principle. Following the reasoning process, the system is capable of taking actions such as searching for information, indexing data in databases, and employing expert AI models. By integrating this information, the system effectively simulates a highly knowledgeable doctor. Working in concert, these agents can deliver precise, explainable solutions to user inquiries.
ClinicalAgent integrates diverse machine learning models and data sources to predict clinical trial duration, failure reasons, outcomes, and enrollment difficulty. As the first multi-agent framework for clinical trials using advanced LLM technology, it features a developed website and aims to become a community platform, providing precise and explainable solutions to user inquiries. In the following paragraph, we will introduce an example of how to use ClinicalAgent to predict clinical trial outcomes.
3.2. Agent Roles and Responsibilities
@@
The ClinicalAgent framework integrates a diverse array of specialized agents, each employing the ReAct and LEAST-TO-MOST reasoning methods to meticulously plan their actions. Through the use of advanced search capabilities, access to specialist models, and indexing in databases, these agents are able to execute a wide range of tasks effectively. Below, we delve into the specific roles and responsibilities assigned to each agent within the system.
3.2.1. Planning Agent
@@
3.2.2. Efficacy Agent
The Efficacy Agent is a specialized module within our multi-agent framework, primarily focused on assessing the therapeutic effectiveness of drugs against specified diseases. This agent utilizes advanced data retrieval and analysis techniques, drawing from rich biomedical databases such as DrugBank (Wishart et al., 2018) and the HetioNet Knowledge Graph to ensure comprehensive and accurate evaluations.
@@ 3.2.2. Efficacy Agent
3.2.3. Safety Agent
The Safety Agent is integral to our ClinicalAgent framework, focusing specifically on the assessment of drug safety and its implications for patient health. This agent leverages a comprehensive repository of pharmacological data and historical clinical trial outcomes to evaluate the risks associated with specific drug-disease interactions. Utilizing databases such as DrugBank and clinical trial registries, the Safety Agent provides detailed insights into the historical safety profiles of drugs.
@@ 3.2.3. Safety Agent
3.2.4. Enrollment Agent
Proper enrollment ensures that the trial has enough participants to statistically power the study. This is essential to detect the true effect of the intervention being tested. Insufficient enrollment can lead to inconclusive or unreliable results because the sample size determines the ability of a trial to accurately reflect the effects of a treatment. We design a Hierarchical transformer-based model that takes eligibility criteria as an input feature and predicts the success rate of enrollment. It is a binary classification problem, where 1 denotes the successful enrollment while 0 does not. The details of the model can be found in Section A.1 in the Appendix.
3.3. Calling External Tools
@@
Drug Databases: Drugbank (Wishart et al., 2018) stands out as a premier resource, offering detailed drug data, including chemical, pharmacological, and pharmaceutical information, with a focus on comprehensive drug-target interactions. Drugbank is not only a repository of drug information but also serves as an invaluable tool for bioinformatics and cheminformatics research, providing data for over 13,000 drug entries including FDA-approved small molecule drugs, FDA-approved biopharmaceuticals (proteins, peptides, vaccines, and allergenics), and nutraceuticals.
Knowledge Graphs: Hetionet (Himmelstein et al., 2017) is an integrative network of biology that encompasses a comprehensive collection of biological entities and their relationships. It uniquely combines data from various biomedical databases covering diseases, genes, compounds, and more, into a single, coherent graph structure. This interconnected approach allows for multifaceted analyses, including drug repurposing, genetic associations, and network medicine. Hetionet includes over 47,000 nodes of different types (e.g., diseases, drugs, genes) and more than 2 million relationships, offering a rich dataset for computational biology and drug discovery.
@@
Enrollment Model: The enrollment model is designed to predict the likelihood of successful participant enrollment in clinical trials based on the eligibility criteria, the drugs involved, and the diseases targeted (Yue et al., 2024b). This is a hierarchical transformer-based model, integrating sentence embeddings from BioBERT (Lee et al., 2020) to capture the nuanced medical semantics in the criteria text. In practice, the Enrollment Agent receives a query containing the drugs, diseases, and detailed eligibility criteria. It processes this information to predict the enrollment difficulty, which aids in planning and adjusting recruitment strategies for clinical trials. This capability supports more efficient trial design and can significantly impact the speed and success of new drug developments. The details can be found in Section A.1 in Appendix.
@@
3.4. Integration of Reasoning Technology
To further enhance the agent’s decision-making capabilities, we integrate advanced reasoning technologies such as ReAct (recognition, action, and context) (Yao et al., 2023) and the Least-to-Most reasoning framework (Zhou et al., 2023). These methodologies complement each other by providing robust mechanisms for addressing complex problems through structured and contextual analysis.
@@ 3.4. Integration of Reasoning Technology
3.5. Workflow
The workflow of our ClinicalAgent system is designed to optimize the collaboration and efficiency of multiple specialized agents to address complex medical inquiries. The process is structured in several sequential steps, as described below:
@@ 3.5. Workflow
This structured workflow ensures that ClinicalAgent effectively mimics a collaborative team of medical specialists, offering precise and comprehensive solutions to complex medical inquiries.
4. Experiment
This section outlines the experimental design used to assess the performance of ClinicalAgent in the setting of clinical trials. We aim to demonstrate that the ClinicalAgent framework significantly outperforms direct predictions by LLM. Additionally, compared to traditional machine learning models, ClinicalAgent also shows competitive results. It is important to note that all machine learning models can be integrated as part of ClinicalAgent; this comparison between end-to-end ML models is merely to showcase the potential of ClinicalAgent.
4.1. Baseline Methods
@@
This comparative analysis will help in highlighting the strengths and potential areas for improvement in ClinicalAgent, guiding future enhancements in the model’s architecture and its application in clinical trial settings.
4.2. Experimental Setup
Our experimental framework was implemented on a server equipped with an AMD Ryzen 9 3950X CPU, 64GB RAM, and an NVIDIA RTX 3080 Ti GPU. We utilized Python 3.8 for scripting and PyTorch for model implementation and training. For each experiment, we used the same seed to ensure reproducibility. For our experimental validation, we randomly selected 40 training samples from the clinical trial outcome prediction benchmark provided in (Fu et al., 2022, 2023). We applied the same approach to select 40 samples from the test set. The decision to use 40 samples was driven by the high computational cost associated with calling the OpenAI API, which necessitated a balance between thorough testing and resource constraints. By ensuring that our training and testing datasets were balanced and randomized, we aimed to provide a robust evaluation framework for our model’s performance while managing the associated costs effectively.